### v2

# Vib2Mol: from vibrational spectra to molecular structures—a versatile deep learning model

Abstract
There will be a paradigm shift in chemical and biological research, to be enabled by autonomous, closed-loop, real-time self-directed decision-making experimentation. Spectrum-to-structure correlation, which is to elucidate molecular structures with spectral information, is the core step in understanding the experimental results and to close the loop. However, current approaches usually divide the task into either database–dependent retrieval and database–independent generation and neglect the inherent complementarity between them. In this study, we proposed Vib2Mol, a general deep learning model designed to flexibly handle diverse spectrum-to-structure tasks according to the available prior knowledge by bridging the retrieval and generation. It achieves state-of-the-art performance, even for the most demanding Raman spectra, over previous models in predicting reaction products and sequencing peptides as well as analyzing experimental spectra and integrating multi-modal spectral data. Vib2Mol enables vibrational spectroscopy a real-time guide for autonomous scientific discovery workflows.111codes are available at https://github.com/X1nyuLu/vib2mol
keywords:
deep learning, vibrational spectroscopy, spectrum-to-structure
1 Introduction
With the rapid development of automated experimental design and execution[1, 2], it has become possible to explore potential chemical reactions and study complex life processes with a closed-loop workflow without human intervention. It may significantly accelerate material design and drug discovery. The key to automating such a closed-loop workflow is to design and execute the next experiment on the basis of the prior knowledge. However, this is particularly challenging due to the lack of quantification for the merits of the decisions. In this context, spectra, especially those obtained from in-situ measurements, have become the key to addressing this challenge by providing the basic structural information of molecules thus offering feedback for each decision. Therefore, it is urgent to develop efficient methods to elucidate molecular structures on the basis of spectral information, i.e., spectrum-to-structure correlation.
Leveraging its superior ability to process big data and uncover latent patterns, deep learning (DL) has significantly advanced the spectrum-to-structure tasks. These DL-based methods can be generally categorized into two ways: database–dependent retrieval and database–independent generation. Retrieval-based approaches, including spectrum-spectrum and spectrum-structure retrieval, rely on comparing the to-be-determined spectrum with candidate spectra or molecular structures according to certain rules to find the best match. These approaches are effective in identifying chemicals within the library that has been previously established or delineated on the basis of prior knowledge, such as DeepSearch[3], FastEI[4] and CReSS[5]. However, these methods inevitably face severe limitations when dealing with out-of-library compounds, owing to the big gap between available experimental spectrum-structure pairs ($\sim 10^{6}$) and vast chemical space[6]. In contrast, generation-based approaches, including conditional generation and de novo generation, seek to predict molecular structures directly from spectra, bypassing the establishment and retrieval of databases. These approaches have shown great promise for predicting previously unidentified chemicals[7, 8, 9, 10, 11, 12, 13, 14]. However, the spectral signal obtained from single technique unveils only a partial view of molecular structure. As a result, the process of converting one type of spectral data into its molecular structure is inherently challenging, let alone the complexity and noise in the experimental spectrum.
Indeed, retrieval is efficient enough to determine in-library molecules, whereas generation becomes the only option for interpreting spectra of out-of-library molecules. However, up to now most of the existing methods have either retrieval or generation but not both. Such a paradigm not only makes model unable to provide appropriate solutions as prior knowledge and databases change, but also ignores the synergy between retrieval-based and generation-based spectrum-to-structure tasks, while this synergy could further improve the performance of spectral annotation. As a result, it is ideal to develop a general model that is capable of retrieval and generation simultaneously and provides dynamic solutions on the basis of available knowledge and databases.
In this study, we propose a DL-based vibrational spectrum-to-molecular structure model (Vib2Mol) to flexibly address a variety of spectral annotation tasks according to the available prior knowledge. We focus particularly on Raman spectroscopy, a widely used non-invasive, in-situ method. Raman spectroscopy has long faced challenges in accurately simulating the Raman spectra of target species, which requires very demanding calculation of high-order energy and force derivative. It becomes even challenging to achieve satisfactory alignment between theoretical and experimental spectra and interpret the experimental spectra. If we can demonstrate the feasibility of Vib2Mol for such intricate form of spectra, if may be easily extended to other spectroscopic techniques. In specific, Vib2Mol adopts an encoder-decoder transformer architecture, and is trained with the strategy of multi-task learning and integrates a wide variety of spectrum-to-structure tasks into one versatile model. Our model achieves state-of-the-art performance on 9 out of 10 test sets when compared to mainstream DL-based methods. It further enhances the accuracy in interpreting spectra of reaction products and peptide sequencing as more knowledge of target molecules is introduced. Additionally, it demonstrates robustness in analyzing experimental spectra and potential for integrating multi-modal spectra from multiple techniques. This advancement demonstrates significant potential for in-situ intelligent analysis of dynamic chemical transformations and biological processes.
2 Results
2.1 Multi-task learning framework: correlating vibrational spectrum and molecular structure
The workflow of Vib2Mol during pre-training, including alignment and generation modules is illustrated in Figure 1. The alignment module (Figure 1A) aims to bring the spectral and structural features of the same molecule as close as possible while separating the features of different molecules simultaneously. Spectra and molecular structures are represented as patch tokens and SMILES tokens, and then encoded into spectral and molecular embeddings by encoders, respectively. These two embeddings are effectively aligned through contrastive learning (CL), enabling cross-modal spectrum-structure retrieval. Since spectrum-spectrum retrieval is implicitly included in the spectrum-structure contrastive learning, there is no need to explicitly design a specific loss function dedicated to spectrum-spectrum retrieval only.
Figure 1B depicts the workflow of conditional generation and de novo generation of molecular structures. Conditional generation, i.e., predicting the occluded molecular structure on the basis of the spectrum, draws on masked language modeling (MLM). Briefly, SMILES tokens, representing the molecular structure, are randomly masked by 45% and then processed by molecular encoders to generate molecular features. Note that molecular encoders share parameters with the alignment module in Figure 1A. Then molecular decoders fused information from both masked molecular embeddings and spectral features, and predicted the to-be-determined tokens using cross-attention. Differently, De novo generation draws on language modeling (LM). Briefly, SMILES tokens are sequentially masked from left to right and directly input into the molecular decoders sharing parameters with MLM. Guided by spectral features and previously generated SMILES sequences, the decoders can predict the next SMILES token from left to right until the entire sequence is complete.
Figures 1C to 1F illustrate the workflow of Vib2Mol during application and inference. (1) For spectrum-spectrum retrieval (Figure 1C), instead of directly comparing spectral similarity by metrics such as Pearson correlation coefficient, the to-be-determined spectrum is encoded into an embedding vector, and the cosine similarity is calculated between this vector and the known spectral embedding vector in the database. (2) For spectrum-structure retrieval task (Figure 1D), Vib2Mol uses CL to minimize the distance between the spectral and structural features of the same molecule, referring to the mainstream spectrum-structure contrastive learning models[3, 5, 15]. spectrum-structure retrieval is based on spectrum-structure similarity, i.e., the cosine value between the spectral embedding of the to-be-determined spectrum and the molecular embeddings of known molecules in the database. (3) For conditional generation task (Figure 1E), Vib2Mol adopts the encoder-decoder architecture. Both the spectrum and partially masked molecular structure are encoded and then fused through the molecular decoder to generate the SMILES of the masked part. (4) For de novo generation (Figure 1F), Vib2Mol directly employs molecular decoders to predict SMILES one by one on the basis of the encoded spectral features until a complete molecular structure is generated. When generating the next token, stochastic perturbation beam search (BS, see Methods for details) is used to ensure the diversity of the results, which is also confirmed to improve the generation performance by ablation experiments (Table S1).
It is worth noting that Vib2Mol adopts staged pre-training (SPT), the advantage of which is confirmed by ablation experiments in Table S1. In the first stage, the alignment module (including the spectral and molecular encoders) is trained by CL loss. After that, the parameters of the trained spectral and molecular encoders are frozen. In the second stage, the generation module (molecular decoder) is trained by MLM and LM losses. The molecular encoder of Vib2Mol shares parameters for CL and MLM, and the molecular decoder shares parameters for MLM and LM. Meanwhile, the spectral features extracted by the spectral encoder are reused all the time. It is the parameter sharing and feature reusing strategies that enables Vib2Mol to address above spectrum-to-structure tasks by simply changing the combination of encoders and decoders, without the need for additional fine-tuning or training.
Figure 1: (A-B) The architecture of Vib2Mol for pretraining and (C-F) the workflow for different spectrum-to-structure tasks. (A) The alignment module: spectra and molecular structures are represented as patch tokens and SMILES tokens, respectively. After processed by their encoders, spectral and molecular information are aligned by CL. (B) The generation module: for conditional generation, molecules are randomly masked 45% and encoded by the same molecular encoder used for spectrum-structure alignment. The molecular decoder fuses spectral information with molecular features and predicts masked tokens. MLM loss is used to compare the distance between predicted and original SMILES. For de novo generation, molecule is sequentially masked and directed input into the same molecular decoder as conditional generation without the prior encoding. Then, the decoder predicts the next token based on previous information and spectral features. By flexibly combining encoders and decoders, Vib2Mol can address four spectrum-to-structure tasks at the same time: (C) spectrum-spectrum retrieval, where only the spectral encoder is used to calculate the similarity between spectral pairs; (D) spectrum-structure retrieval, where spectra and molecules are encoded by their respective encoders to determine spectrum-structure similarity; (E) conditional generation, and (F) de novo generation, both following workflows during the stage of pretraining.
2.2 State-of-the-art performance of Vib2Mol
Figure 2A shows the performance of Vib2Mol on the vibrational spectrum-to-structure benchmark (ViBench, see Methods for details), in comparison with three mainstream DL-based models: vanilla-CL, vanilla-MLM, and vanilla-LM (see methods for more details about baseline models). Vib2Mol is not only capable of addressing all tasks simultaneously but also achieves the best performance on 9 of 10 test sets (detailed metrics are recorded in Table S1).
For spectrum-spectrum retrieval, the Recall@1 of Vib2Mol (76.46%) in the VB-geometry is comparable to that of vanilla-CL (76.90%), both of which are significantly better than traditional methods (Recall@1 $\leq$ 34.71%). Such huge difference is a result of different matching strategies used for these three methods. Instead of directly using the cosine value or Pearson correlation coefficient to evaluate the similarity of spectral pairs, employing deep learning to project spectrum into the latent space is important. For spectrum-structure retrieval, on the VB-mols test sets, Vib2Mol can further increase the Recall@1 of vanilla-CL from 78.07% to 78.92%.
For the generation tasks, Vib2Mol can not only address them simultaneously but also achieve better performance. Taking the VB-mols as an example, the molecular accuracy of Vib2Mol, a metric for conditional generation (see more in Methods), reached 92.54%, better than vanilla-MLM (90.67%), and the Recall@1 of de novo generation was 56.16%, better than that of vanilla-LM (55.15%).
It is worth noting that the performance of all models on VB-qm9 is generally better than that on VB-zinc15. This is due to the much less molecular diversity of VB-qm9 than that of VB-zinc15, resulting in a smaller search space in retrieval and generation. Overall, the Vib2Mol model outperforms dedicated models designed for single tasks, indicating a synergistic effect between multiple spectrum-to-structure tasks.
2.3 Synergistic interaction among modules of Vib2Mol
The synergistic interaction was explored by recursively combining three single-task models, including spectrum-structure retrieval model trained with CL only, conditional generation model trained with MLM only, and de novo generation model trained with LM only. As shown in Figure 2B and Table S2, on the test set of VB-mols, the best model for spectrum-structure retrieval is CL-only. However, as more tasks are introduced (CL→CL+MLM→CL+MLM+LM), the retrieval performance gradually decreases (Recall@1 drops from 78.92% to 76.26% and 75.35%). In contrast, the performance in de novo generation continuously improves (Figure 2D) with the Recall@1 increasing from 54.99% (LM-only) to 57.84% (CL+LM) and then 58.68% (CL+MLM+LM). Meanwhile, for conditional generation, the accuracy increases from 93.59% to 94.03% with the introduction of CL, then decreases to 92.85% with the further introduction of LM (Figure 2C).
To further investigate the mechanism of the interaction among three tasks, taking CL+MLM+LM as an example, the relative contribution of each loss to updating parameters was quantified (Figure S1, see Methods for the detailed calculation). With the increase of epochs, the contribution of CL loss significantly decreases and approximately equals to 0% when epoch is larger than 200. In contrast, the contribution of MLM loss remains around 20-40%, and that of LM loss always dominates while gradually increasing.
From the perspective of difficulty of task, generation requires searching for molecular structures that meet multiple constraints in an open chemical space[16], while the smaller search space and more clear problem definition both offer a lower complexity of retrieval[17]. As a result, generation losses (MLM and LM) are more likely to dominate the optimization[18, 19, 20], resulting in weaker retrieval performance. At this point, CL loss with lower contribution acts similarly to auxiliary regularization and leads to stronger generation performance. Interestingly, the conditional generation metric of CL+MLM falls between MLM-only and CL+MLM+LM due to the moderate contribution of MLM. This phenomenon further confirms the order of these tasks: spectrum-structure retrieval (normal),conditional generation (harder), and de novo generation (hardest). In summary, the simpler task has lower loss contribution, promoting the performance of harder task with higher loss contribution.
Figure 2: (A) Performance of single-task models and Vib2Mol (red region) on 10 test sets. Blue, purple, and green regions represent retrieval, conditional generation, or de novo generation tasks, respectively. Ablation experiments of multi-task learning for (B) retrieval, (C) conditional generation, and (D) de novo generation.
2.4 Visualizing molecular and spectral embeddings to elucidate compound class clusters
In order to verify whether the superior performance of Vib2Mol stems from its unbiased understanding of the semantic information of molecules and spectra, the following two experiments were carried out.
From the macroscopic perspective, we randomly selected one million molecules from VB-qm9 and calculated their structural and spectral similarities. As shown in Figure 3A, on the one hand, a significant positive correlation trend (Pearson correlation coefficient is 0.65) is between the molecular similarity and spectral similarity, both of which are extracted by Vib2Mol. Such correlation indicates that the model can well align the structural and spectral information of molecules. On the other hand, there is an obvious variance between the two similarities during the range of 0-0.5, indicating that there are certain differences between them. We thereby sampled spectrum-molecule pairs containing common functional groups (ether, alcohol, amine, nitrile, amide, ketone, alkene, haloalkane, alkyne, imide). Subsequently, t-SNE was employed to visualize the distribution of structural and spectral embeddings of molecules with different functional groups. Molecules with the same functional group can be clustered to several groups either by structural embeddings (Figure 3B) or spectral embeddings (Figure 3C). In addition, the overall distribution and clustering of structural or spectral embeddings are similar, confirming the reliable performance of Vib2Mol in aligning molecular structures and spectra.
From the microscopic perspective, although similar molecular embeddings are close to each other in distance, molecules with the same functional group often gather into several clusters, due to the fact that molecular structures consist of both functional groups and backbones. For instance, although molecules 1 and 2 are classified as alcohol, the difference in backbones, in which a carbon ring is for molecule 1 and a heterocycle containing O and N is for molecule 2 (Figure 3D), brings very low molecular similarity (0.02, Figure 3E) and spectral similarity (0.23, Figure 3F) between these two molecules, respectively. In contrast, although molecule 2 and molecule 3 (nitrile) are grouped to different clusters, the much similar backbone (both are heterocycles containing O and N) offers both higher molecular similarity (0.32) and spectral similarity (0.30) between them than those between molecules 1 and 2.
In summary, at the macro level, Vib2Mol can well understand the structural and spectral information of molecules and effectively extract the relationship between them. At the micro level, a tight connection is between molecular similarity and spectral similarity, but with a certain gap. This phenomenon is not surprising at all, since Raman spectrum cannot completely reveal all the structural information, but only reflect part of the characteristics of a molecule from the aspect of vibration, resulting in a deviation between the calculated similarity and actual molecular characteristics.
Figure 3: (A) Joint distribution of molecular similarity and spectral similarity. Visualization of (B) molecular and (C) spectral embeddings using t-SNE. (D) Molecular structure and related Raman spectra of selected molecules. Similarity matrices of molecular (E) and spectral (F) embeddings.
2.5 Generating products for chemical reaction
The autonomous robot laboratory is leading new paradigm shifts in fields such as chemical synthesis, catalysis and drug screening[1]. The related advancements are underpinned by spectral information provided by various techniques. For instance, the autonomous and efficient exploration of chemical synthesis (such as combinatorial small-molecule synthesis, designing supramolecular materials, and screening photocatalysts) can be achieved with the aid of HPLC-MS and NMR[2]. However, applying current spectrum-to-structure methods to real conditions remains challenging. On the one hand, researchers have different levels of knowledge about different synthetic methods. As a result, it is crucial to fully utilize the available prior knowledge to help select appropriate molecular elucidation strategies. On the other hand, spectra measured in practice are often mixtures of reactants and products. It is a key issue to elucidate molecules under the interference of impurities. The solvation of these two problems by Vib2Mol were demonstrated as follows.
Taking the product prediction in substitution reaction of polycyclic aromatic hydrocarbons (PAHs) based on Raman spectroscopy as an example (see Methods for details of dataset), there may have three situations.
(1) Spectrum-structure retrieval is for the well-known reactions, i.e., predicting the specific substitution site with the known type of substituent. Due to the limited substitution sites of PAHs, it is possible to retrieve by traversing all possible substitution structures, thereby outputting the structure with the highest spectrum-structure similarity. As shown in Table 1, the Recall@1 of Vib2Mol for benzene, naphthalene, and anthracene reached 99.57%, 99.75%, and 99.14%, respectively, indicating Vib2Mol can nearly perfectly perform spectrum-structure retrieval within the limited search space. Obviously, as prior knowledge decreases, the potential research space significantly increases, making it difficult to traverse all possible structures, and the generation is highly demanded.
(2) Conditional generation is for the partial known reactions, i.e., predicting the type of substituent with the known substitution. The average Recall@1 of Vib2Mol is 98.95% in predicting one unknown substituent, and remains at 96.64% for the case of two unknown substituents. Note that the accuracies here are somewhat inflated. Before performing conditional generation, it is necessary to design certain “blanks” for the model to “fill in”, but Vib2Mol directly replaces the characters at the corresponding positions with “<mask>”. This approach may leak the number of characters to be filled. Although we tried to change the number of characters corresponding to “<mask>”, it is hardly to exhaust all possibilities. Obviously, this shortcoming should to be addressed in the future.
(3) De novo generation is for the new reactions, i.e., predicting a completely unknown molecular structure, including the type of substituents and all substitution sites, simultaneously. Due to the simplicity of the structure, the Recall@1 of benzene (95.32%) is significantly better than that of naphthalene (82.44%) and anthracene (84.91%). The average Recall@1 of the three situations can reach 86.63%. Such a high Recall@1 is mainly because of the limited search space of this case, let alone the real condition predicting by a mixed Raman spectrum.
Therefore, by extracting approximately 15,000 chemical reactions listed in the second World AI4S Prize-Material Science Track[21], we calculated the Raman spectra of reactants and products, and mixed them according to the labeled yields (ignoring the differences in Raman scattering cross-sections of different molecules) to simulate the mixed spectra measured in real condition (see Methods for details).
When the expected product is in the database (Figure S2A), Vib2Mol trained on unmixed spectra (Vib2Mol-unmix) achieved a Recall@1 of 84.87% for spectrum-to-structure retrieval on the unmixed spectrum test set. However, performance dropped significantly to 40.47% on the mixed spectra test set, which is closer to the real-world condition. In contrast, Vib2Mol trained on mixed spectra based on yields (Vib2Mol-mix) achieved a Recall@1 of 82.70%. When the expected product is out of the database (Figure S2B), Vib2Mol-unmix achieved a Recall@1 of only 10.17% for de novo generation on the mixed spectrum test set, while Vib2Mol-mix achieved a Recall@1 of 24.40%. These results clearly demonstrate that introducing yield information significantly enhances Vib2Mol’s ability to annotate mixed Raman spectra.
Considering the uncertainty of yields in real experiments, we replaced the fixed yield with a random value within a certain range and further investigated the effect of the range of yield values on performance (Figure S2). The comparable Recall@1 to that with labeled yields implies the retrieval performance is not sensitive to the sampling range of yields. Differently, thanks to the increased diversity of training data with the expanded sampling range, this strategy improves the performance of generation.
It is worth noting that the model trained by unmixed spectra has a Recall@1 of 25.55% for de novo generation on the unmixed spectrum test set, while the model trained with random yields of 10-100% achieved a Recall@1 of 27.65% on the mixed spectrum test set. Such comparison indicates that mixing spectra can slightly enhance generation capabilities of Vib2Mol. The reason for this phenomenon is currently unclear. One possibility is the strengthened characteristic peaks by mixing Raman spectra of reactants (Figure S3), thereby assisting in the generation of product structures.
2.6 Peptide sequencing and PTMs identification
Native proteins are composed of 20 amino acids and their post-translational modifications (PTMs). As sequences determine the structures and functions of proteins, protein sequencing and the identification of PTMs sites are key issues in reveals the functions and mechanisms of proteins in cellular function regulation, gene expression regulation, signal transduction, and the occurrence and development of diseases. To simplify the complexity of protein sequences, a bottom-up strategy is commonly adopted, which involves generating peptides of varying lengths (1-4 amino acids or longer) through chemical or enzymatic cleavage. By sequentially identifying these peptides, de novo sequencing can be achieved. However, considering the vast sequence space of polypeptides ($20^{N}_{AA}$), efficiently identifying the 20 amino acids and their combinations remains highly challenging[22]. Although the unique fingerprint vibrational information in Raman spectrum is for each biomolecule (i.e., DNA, proteins)[23, 24, 25], the complexity of Raman spectra of peptides hindered the systematic identification of polypeptides for de novo protein sequencing. Limiting the length of peptide sequences to tetrapeptides or shorter, we tried to infer peptide sequence using its Raman spectra by Vib2Mol.
The Vib2Mol pre-trained on VB-mols was fine-tuned by peptides represented by SMILES. As shown in Table 1, the Recall@1 of the model (Vib2Mol-SMILES) for spectrum-structure retrieval is 63.11%, when the to-be-determined peptide is in the database. Otherwise, the Recall@1 for de novo generation drops to 27.55%. The low Recall@1 ignited us to change peptide representation from SMILES to residue sequences, considering the relatively patterned residue structure. The obtained Vib2Mol-sequence model significantly improved Recall@1 for retrieval and generation up to 67.58% and 39.92%, respectively. This improvement is mainly because of the drastically reduced token length by residue sequences, thereby reducing the complexity and improving the accuracy of sequence generation. This is the reason why current models[3, 12, 26] are mainly based on residue sequences. It is not surprising that the Vib2Mol sequence performed best in elucidating dipeptides (94.74% and 89.47% for retrieval and generation, respectively). Since the search space increases exponentially with the number of residues, the performance of Vib2Mol-sequence gradually decreases with the increasing length of peptides. Nevertheless, even for tetrapeptides, the Recall@1 for retrieval and generation can still reach 65.48% and 36.62%, respectively.
As for the identification of PTMs sites, we constructed the VB-peptide-mod dataset, which includes the most representative phosphorylation and sulfation (see Methods for details), and fine-tuned Vib2Mol-sequence by it. Depending on the level of prior knowledge in practical application, there are three cases (Table 2).
(1) Determining modification type at a specific residue site. Vib2Mol can achieve a high accuracy (75.03%) for the three categories (sulfation, phosphorylation, and unmodified) by conditional generation. (2) Retrieval of peptides within the database. Vib2Mol can reach a Recall@1 of 65.03% by calculating spectrum-structure similarities. (3) De novo sequencing for peptides outside the database. A Recall@1 of 27.99% can be achieved by generation module. As a proof-of-concept, Vib2Mol demonstrates the feasibility of using theoretical Raman spectra in de novo sequencing of peptides and identifying PTMs sites. This advancement holds significant promise for applications in biomedicine, immunology, and drug development. We anticipate a synergistic integration with experimental data will further enhance its utility and uncover new insights.
Table 1: Effect of representation and length of peptide on performance.
Peptide Retrieval De novo generation
Recall@1 Recall@3 Recall@1 Recall@3
Vib2Mol-SMILES 63.11 88.13 27.55 33.87
Vib2Mol-sequence 67.58 89.89 39.92 51.18
Dipeptide 94.74 100 89.47 94.74
Tripeptide 77.56 94.45 55.24 68.31
Tetrapeptide 65.48 88.95 36.62 52.51
Table 2: Performance for various PTMs types under different tasks.
site classification Peptide Retrieval De novo generation
Accuracy Recall@1 Recall@3 Recall@1
Unmodified 79.16 76.02 93.05 27.26
Phosphorylated 74.97 62.31 86.54 28.94
Sulfated 71.94 58.71 84.68 27.86
Averaged 75.03 65.03 87.77 27.99
2.7 Generalization of Vib2Mol for experimental spectra and multimodal spectral analysis
The above two applications demonstrate the power of Vib2Mol in spectrum-to-structure tasks based on theoretical Raman spectroscopy from chemistry to biology. As it is well-known that there is a huge gap between theoretical and experimental spectra, due to the difficulty in theoretically simulating the interactions among different vibrational modes. Therefore, we explored the feasibility of Vib2Mol in experimental infrared spectroscopy (NIST, containing about  12,000 experimental infrared spectrum-molecular structure pairs), considering the current lack of a large-scale experimental Raman spectroscopy database for molecular elucidation.
As expected, due to the significant differences in spectral features between theoretical and experimental spectra (Figure S4), the Vib2Mol model trained on theoretical spectra performed poorly on experimental database. As shown in Table 3, the Recall@1 for spectrum-structure retrieval and de novo generation were almost 0%, with only conditional generation achieving an accuracy of about 25%. The certain level of accuracy of conditional generation is possibly due to the contextual information of the unmasked structure. When Vib2Mol was trained on experimental infrared spectra, the Recall@1 for spectrum-structure retrieval increased to 19.73%, the accuracy for conditional generation increased to 48.89%, and the Recall@1 for de novo generation increased to 4.41%. Considering the distance is far away from the requirements of practical applications, we pre-trained Vib2Mol on theoretical database, and then employed experimental infrared spectra to fine-tune, resulting in a significant improved performance. The Recall@1 for spectrum-structure retrieval increased to 51.49%, the accuracy for conditional generation increased to 71.55%, and the Recall@1 for de novo generation increased to 11.79%. This improvement is likely due to a certain degree of commonality between theoretical and experimental spectra. Therefore, the pre-training by theoretical spectra provides some assistance in interpreting experimental spectra. No wondering, the performance could be enhanced by introducing other molecular information, such as the chemical formula[9, 14, 27], which is out of the scope of current discussion.
On the basis of the theoretical infrared and Raman databases, we further examined the impact of integrating multi-modal spectral information on Vib2Mol performance. As shown in Table 4, on the spectrum-structure retrieval of VB-mols, there is no much difference in performance between models trained on single-spectral and dual-spectral information. Among them, Raman-only performs the best, with a Recall@1 of 78.92%, while the performance of utilizing Raman+IR is slightly worse (78.83%). One possible reason is that retrieval aims to focus on clear, unique, and highly discriminative features, in order to accurately find the entry in a large number of known molecular structures that best matches the given spectrum. However, Infrared and Raman spectra, observing molecular vibrations from different perspectives, are not completely aligned and unbiased. Therefore, there may be some inconsistency in certain vibrational features, which may have a negative impact on the integration[28]. To reduce this negative impact, one could consider introducing other spectral information from different aspects, such as NMR and MS, to provide more perspectives for a more comprehensive observation of molecular structure features.
On conditional generation and de novo generation, the synergistic effect of the two spectroscopies is very convincing. Taking the de novo generation on VB-mols as an example, the Recall@1 for Raman-only was 59.51%, and the introduction of infrared spectra increased it by 5.5% to 65.01%. This improvement likely stems from the fact that generation tasks prioritize feature comprehensiveness and correlation over highly discriminative features. The diverse molecular structure information provided by different spectra can offer the model richer clues and a more integrate framework for constructing molecular structures or spectra. Therefore, even if some features are not precise enough, the model can still generate reasonably based on the overall information, resulting in better performance.
Table 3: Performance of models trained with different data set tested on NIST-based experimental infrared spectral dataset.
Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
Theo. IR 0.1 25.61 0
Exp. IR 19.73 48.89 4.41
Theo. IR + Exp. IR
(finetune) 51.49 71.55 11.79
Table 4: Testing performance of models trained with different datasets.
qm9 zinc15 mols qm9 zinc15 mols qm9 zinc15 mols
Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
Theo. IR 80.19 76.02 77.66 97.95 86.23 92.5 57.58 45.16 53.98
Theo. Raman 81.51 76.65 78.92 97.91 87.04 92.88 63.31 49.86 59.41
Theo. Raman + IR 81.03 77.39 78.83 98.35 88.65 93.85 69.38 54.29 65.01
* The “qm9”,“zinc15” and “mols” here all refer to specific subsets within the ViBench.
3 Discussion
In this study, we proposed Vib2Mol, a DL model for vibrational spectroscopy, which can effectively address multiple spectrum-to-structure tasks according to available prior knowledge. On ViBench, Vib2Mol outperformed traditional methods on 9 of 10 test sets. Such outstanding performance stems from the synergistic of retrieval and generation modules which lead to the better establishment of the correlation between spectrum and molecular structure, which was confirmed by effective clustering molecules with similar structures.
Vib2Mol has shown substantial potential in chemical and biological applications, where we have further tackled several unexplored challenges with in-silico data. On the one hand, chemical reactions inevitably lead to a mixture of reactants and products, which thus results in mixed spectra posing a challenging issue for spectral annotation. We showcased the capability of Vib2Mol to interpret mixed-spectra, which achieved the recall@1 of 82.70% and 24.40% for retrieval and de novo generation on chemical reaction dataset with real yields, respectively. On the other hand, Vib2Mol enables Raman spectroscopy as a unique omics method which not only achieved a Recall@1 of 39.9% for de novo peptide sequencing, but also efficiently predicted PTMs sites of phosphorylated and sulfated modification, where the traditional mass spectrometry falls short. After fine-tuned by experimental spectra, Vib2Mol can be adapted to interpret experimental infrared spectra. Finally it demonstrated potential for integrating multi-modal spectra from multiple techniques as well.
Vib2Mol demonstrates the potential for in situ monitoring of dynamic chemical reactions and life processes on the basis of vibrational spectroscopy. In the future, to better elucidate molecular conformations in dynamic processes, a possible improvement lies in the introduction of stereochemical information. In addition, it is also of great interest to design more flexible generative modules to equip the models with bidirectional spectrum-to-structure and structure-to-spectrum predictions.
4 Methods
4.1 Reference data
We have established a vibrational spectrum-to-structure benchmark (ViBench, VB), which consists of eight parts: VB-qm9, VB-zinc15, VB-mols, VB-geometry, VB-PAHs, VB-RXN, VB-peptide, and VB-peptide-mod. Details are listed in supplementary information.
Density functional theory (DFT) was employed to perform conformational optimization of these molecules and calculated the corresponding infrared and Raman spectra. Unless otherwise specified, all quantum chemical calculations were carried out using the Gaussian 09 program[g09]. The geometries were optimized using the B3LYP-D3BJ functional with a 6-311+G** basis set. Frequency calculations were obtained at the same level at the optimized geometry.
Furthermore, to test the generalization of Vib2Mol on experimental spectra, we collected experimentally measured infrared spectra of 12,937 small molecules from the public NIST dataset[linstrom1997nist]. For details on the specific division of the above datasets for training, validation, and testing, please refer to Table S4. To facilitate subsequent calculations, the spectral dimensions were unified to 1024, and molecular structures were all represented using SMILES.
4.2 Related works and baseline models
To fairly compare the performance among Vib2Mol and current methods, we surveyed spectrum-to-structure models based on vibrational spectroscopy in the community. For spectral-structure retrieval, CL is currently the most popular framework. DeepSearch[3], CReSS[5] and SMEN[29] all employ CL to bring the spectra and corresponding structures of the same molecule closer together, achieving great spectral-structure retrieval performance for mass spectrometry, NMR, and IR, respectively. We have borrowed the architecture of the above methods thus building a similar baseline model (vanilla-CL) applicable to Raman spectra. It is worth noting that although SMEN is very similar to Vib2Mol, the two are not comparable because SMEN uses atomic coordinates while Vib2Mol uses SMILES to represent the molecular structure. But in Section 2.3, we have confirmed that the performance of Vib2Mol is better than CL+LM, which is similar to SMEN, in both retrieval and generation.
For the conditional generation, MLM is currently the most popular approach. CO-BERT realized bidirectional prediction between molecular structures and vibrational spectra by MLM[30]. However, CO-BERT focus on predicting atomic coordinates by vibrational spectra and contextual structural information, which cannot be compared with Vib2Mol. Therefore, we followed its architecture and built a similar baseline model (vanilla-MLM).
For the de novo generation, Alberts et al.[14] and Wu et al.[9] both adopted an encoder-decoder architecture and introduced the constraint of molecular formula as well. However, in many cases, information such as the molecular formula is not readily available[7]. Therefore, we followed the above two models and built a formula-free baseline model (vanilla-LM).
4.3 Spectral and molecular representation
As shown in Figure S5A, the convolutional kernels with size of 8 were first used to slice the original spectra into 128 patches. linear projection was then employed to transform each patch into a 768-dimensional vector, i.e., spectral embeddings. As shown in Figure S5B, the preprocessing of molecules is similar. the molecular structure is represented as a SMILES string and is split into several discrete characters, i.e., SMILES tokens. After looking up the codebook, all characters are mapped to 768-dimensional vectors, i.e., molecular embeddings. Subsequently, the <CLS>token, representing the global information of the sequence, was inserted at the beginning of both the spectral sequence and the molecular structure sequence, and positional encoding was added to both. Finally, a 6-layer Transformer encoder based on self-attention was used to update the features of each token in the sequence. It is worth noting that at this point, the spectrum and molecular structure only interact with their own features and do not communicate with each other here.
4.4 Alignment between spectrum and molecular structure
To align the features of spectra and molecular structures, CL was introduced. As shown in Figure S6, spectral and molecular features were extracted from their respective encoders, then a spectrum-structure similarity matrix was obtained through the dot product. By optimizing this matrix, the spectral and molecular embeddings of the same molecule were made as close as possible (with the diagonal elements approaching 1), while the embeddings of mismatched spectrum-molecule pairs were made as distant as possible (with the off-diagonal elements approaching 0).
During the training phase, we used a symmetric cross-entropy loss[31] to calculate the similarity errors between the spectra and structures of the same molecule and updated the neural network based on this. The specific formula of the loss function is as follows:
$L_{total}=\frac{1}{2}(L_{spectrum}+L_{structure})=-\frac{1}{2}(\sum_{i=1}^{m}{ilog(p_{i})}+\sum_{j=1}^{n}{jlog(q_{j})})$ (1)
where $m$ and $n$ are the number of rows and columns of the probability distribution matrix, respectively. $ilog(p_{i})$andj $ljlog(q_{j})$represent the cross-entropy of spectrum-to-structure, and structure-to-spectrum, respectively. During the testing or inference phase, only the dot product of the features of the to-be-determined spectrum and the molecules in the library needs to be calculated, and the top-k results are taken as the final results (Figure S6).
4.5 Spectrum-guided molecular generation
After aligning the spectra and molecular features, we aim to generate molecular structures based on spectral information. During the training phase, we integrated two training tasks, MLM and LM. For MLM, we randomly masked 45% of the content in the structural sequence and utilized cross-attention to enable the model to learn how to restore the masked parts of the structure based on the spectrum and contextual tokens (Figure S5C). For LM, we enforced the model to learn how to predict the next character based on the previously generated text and under the guidance of the spectrum (Figure S5D).
The loss functions for both MLM and LM are based on cross-entropy, as detailed below:
$L_{MLM}=-\frac{1}{N}\sum_{i=1}^{N}{logP(\hat{y}_{i}|y_{unmasked},s)}$ (2)
where $N$ is the total number of masked positions, $y_{unmasked}$ represents the contextual tokens around the masked ones, $\hat{y}_{i}$ represents masked tokens to be predicted, and $s$ is the input spectrum.
$L_{LM}=-\frac{1}{M}\sum_{i=1}^{N}{logP(\hat{z}_{j}|z_{prev},s)}$ (3)
where $M$ is the length of the SMILES sequence, $z_{prev}$ represents the previous generated SMILES, $\hat{z}_{j}$ represents the next to-be-predicted token, and $s$ is the input spectrum.
4.6 Stochastic Perturbation Beam Search
For de novo generation, the greedy search strategy, which only selects the token with the highest probability as the next character may fall into local optima. To enhance the diversity of the generated results, we adopted a beam search method combined with stochastic perturbation. Specifically, first, the log-probabilities ($logP$) of the current candidate tokens were calculated. Then, Gumbel noise was introduced to perturb these log-probabilities, making the selection process of candidate tokens somewhat stochastic, i.e.:
$logP$ $=log(\frac{e^{\frac{p}{\tau}}}{\sum_{j}{e^{\frac{p_{j}}{\tau}}}})$ (4)
$noise$ $=-log(-log(Uniform(0,1)))$ (5)
$score$ $=logP+\alpha\cdot noise$ (6)
where $p$ is the probability distribution of the current token to be predicted, $\tau$ is the temperature of the random perturbation, a larger $\tau$ can make the probability distribution flatter, thereby increasing the randomness of generation, and $\alpha$ is a weighting factor used to balance the contribution of the log-probability and the current noise. In this method, $\alpha$ is set to 0.1.
4.7 Metrics for conditional generation
To better evaluate the performance of conditional generation, we compared two metrics: token accuracy and molecular accuracy. As shown in Figure S7A, token accuracy takes each character to be predicted as the smallest granularity and assesses the model’s ability to restore the masked characters. However, the same molecule can be represented by different SMILES. Therefore, molecular accuracy does not examine the correctness of each character but is designed to evaluate whether the finally predicted molecule is correct (Figure S7B). In addition, we only masked the content between “(” and “)”, so as to ensure that all parts to be predicted are complete branch structures which have clear structural information rather than random combination of characters.
4.8 Calculating contributions of different losses
Deep learning models calculate the gradients of the loss function with respect to the weights through chain rule, thereby updating the weights of the entire network. As shown in Figure S5, the CLS token of the spectrum in Vib2Mol is a learnable vector. Therefore, the gradients of each loss function on this vector can represent their corresponding optimization directions, thus quantifying the contribution of each loss function to the weight update. To further compare the differences between the optimization directions guided by each loss function and the total optimization direction, we calculated the cosine similarity between each gradient and the total gradient and normalized it to the range of 0-1, i.e.:
$\nabla L_{i}$ $=\frac{\partial{L_{i}}}{\partial{t}}$ (7)
$CosSim(\nabla L_{i},\nabla L_{total})$ $=\frac{\nabla L_{i}\cdot\nabla L_{total}}{\parallel\nabla L_{i}\parallel\parallel\nabla L_{total}\parallel}$ (8)
$Contribution_{L_{i}}$ $=\frac{CosSim(\nabla L_{i},\nabla L_{total})}{\sum_{j}{CosSim(\nabla L_{j},\nabla L_{total})}}$ (9)
where $L_{i}$ represents each loss fuction, $t$ represents spectral CLS TOKEN, which is a vector.
[…]
Supplementary information Details about reference data, extra figures and tables are available in the supplementary information.
[…]
Acknowledgements This work was supported by the National Natural Science Foundation (Grant No: 22227802, 22021001, 22474117, 22272139) of China and the Fundamental Research Funds for the Central Universities (20720220009) and Shanghai Innovation Institute.
Appendix A Reference data
We have established a vibrational spectrum-to-structure benchmark (ViBench, VB). As shown in Table S4, the molecular data of VibBench consists of eight parts:
VB-qm9: 133,434 organic small molecules extracted from QM9, composed of C, H, O, N, and F atoms, with the number of heavy atoms less than 10. Each molecule in this subset has only one stable conformation.
VB-zinc15: 50,114 drug molecules extracted from ZINC15[sterling2015zinc], involving a wider range of elements, including C, H, O, N, S, F, Cl, Br, P, and Si, with the number of heavy atoms ranging from 4 to 45. Notably, since the ZINC15 dataset contains many isomers, and VB-zinc15 only ensures the uniqueness of ZINC-IDs, 7,556 molecules in this subset have multiple stable conformations.
VB-mols: For convenience in pre-training and evaluation, we merged VB-qm9 and VB-zinc15, and the combined dataset is referred to as VB-mols. In other words, VB-mols is not an additional dataset but an integration of existing data.
VB-geometry: 7,227 organic small molecules extracted from GEOM, each with two stable conformations. We randomly used the spectrum of one conformation as the query input and the other as the reference spectrum, thus constructing a test set for evaluating the model’s spectrum-to-spectrum matching performance.
VB-PAHs: Includes 1,268 benzene derivatives, 1,853 naphthalene derivatives, and 1,175 anthracene derivatives. The substitution sites for benzene include (1,2), (1,3), and (1,4); for naphthalene, they include (1,2), (1,5), (1,8), (2,6), and (2,7); and for anthracene, they include (1,2), (2,3), and (2,6). All derivatives contain two common substituents as detailed in Table S5.
VB-RXN: 15,639 unique reaction data extracted from The second World AI4S Prize-Material Science Track. Each data entry includes the yield, structures, and Raman spectra of reactant 1, reactant 2, and the product. All molecules have a maximum of 20 heavy atoms and only contain C, H, N, O, F, S, Cl, P, and Br elements.
VB-peptide: Includes 273 dipeptides (68.25% of all possible dipeptides), 4,058 tripeptides, and 21,624 tetrapeptides. All peptides are generated based on the permutations and combinations of A, N, D, C, Q, E, G, H, I, L, M, F, P, S, T, Y, and V.
VB-peptide-mod: Includes 3,815 unmodified peptides, 3,716 phosphorylated peptides, and 5,023 sulfated peptides. All peptides are either tripeptides or tetrapeptides with at most one modification site. The specific modification sites include O-phosphorylation and O-sulfation of tyrosine, serine, and threonine, as well as two different N-phosphorylation modifications of histidine.
Appendix B Figures
Figure S4: Normalized cosine values of the gradients with CL, MLM and LM losses relative to the total gradient at different epochs.
Figure S5: Performance of Vib2Mol trained with different mix ratios on unmixed and mixed spectra.
Figure S6: Examples of (A) substitution and (B) Diels-Alder reactions. (C-D) Raman spectra of reactants, product and mixture for each reaction. Compared to the spectrum of the product, the spectrum of the mixture has more characteristic peaks.
Figure S7: Three examples of comparison between theoretical and experimental spectra sampled from NIST.
Figure S8: Architectures for (A) spectral encoding, (B) molecular encoding, (C) masked language modeling and (D) language modeling.
Figure S9: Workflow of contrastive learning for (A) training and (B) testing.
Figure S10: Comparison between token accuracy and molecular accuracy.
Appendix C Tables
Table S5: Add caption
VB-geometry VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols
Spectral retrieval Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
Cosine similarity 34.67 / / / / / / / / /
$\rho$ similarity 34.71 / / / / / / / / /
vanilla-CL 76.9 80.94 75.19 78.07 / / / / / /
vanilla-MLM / / / / 97.22 83.06 90.67 / / /
vanilla-LM / / / / / / / 58.65 46.05 54.99
Vib2Mol 76.46 81.51 76.65 78.92 97.91 87.04 92.88 63.31 49.86 59.41
Vib2Mol
(w/o RR) 76.46 81.51 76.65 78.92 97.91 87.04 92.88 60.63 47.68 56.88
Vib2Mol
(w/o RR and SPT) 72.98 77.38 74.29 75.35 97.14 87.92 92.85 62.89 48.34 58.68
Table S6: Performance of 6 methods at 4 tasks during the ablation experiments.
VB-geometry VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols VB-qm9 VB-zinc15 VB-mols
Spectral retrieval Spectrum-structure retrieval Conditional generation De novo generation
(Recall@1 / %) (Recall@1 / %) (molecular accuracy / %) (Recall@1 / %)
CL-only 76.46 81.51 76.65 78.92 / / / / / /
MLM-only / / / / 97.75 88.74 93.59 / / /
LM-only / / / / / / / 58.65 46.05 54.99
CL+MLM 75.63 78.33 75.53 76.26 98.27 89.13 94.03 / / /
CL+LM 73.39 72.67 70.07 70.66 / / / 61.89 47.91 57.84
CL+MLM+LM 72.98 77.38 74.29 75.35 97.14 87.92 92.85 62.89 48.34 58.68
Table S7: Performance of Vib2Mol in PAHs.
Categories/ Benzene Naphthalene Anthracene Averaged
Recall@1/
Tasks
Substitution site retrieval 99.57 99.75 99.14 99.54
(given substituents)
Substituent generation 99.57 98.72 98.71 98.95
(partially unknown)
Substituent generation 98.28 96.55 95.07 96.64
(both unknown)
De novo generation 95.32 82.44 84.91 86.63
(totally unknown)
Table S8: Details about the data split for training, validation, and testing.
Datasets # Training samples # Evaluation samples # Testing samples
qm9 93403 13344 26687
zinc15 38089 5442 10883
geometry 0 0 11318
PAHs 3006 430 860
RXN 10947 1564 3128
Peptide 18168 2596 5191
Peptide-mod 8787 1256 2511
NIST-IR (Exp.) 9055 1046 2043
Table S9: All substituents existing in VB-PAHs.
Name Formula SMILES
Trichloromethyl $CCl_{3}$ C(Cl)(Cl)Cl
Dichloromethyl $CHCl_{2}$ C(Cl)Cl
Chloride $Cl^{-}$ Cl
Chloromethyl $CH_{2}Cl$ CCl
Thiyl (Sulfur) $SH$ S
Trifluoromethyl $CF_{3}$ C(F)(F)F
Difluoromethyl $CHF_{2}$ C(F)F
Fluoride $F^{-}$ F
Fluoromethyl $CH_{2}F$ CF
Hydroxyl (Hydroxy) $OH^{-}$ O
Sulfonic Acid $-OSO_{3}H$ S(=O)(=O)O
Phosphonic Acid $PO(OH)_{2}$ P(=O)(O)O
Nitro $-NO_{2}$ N+[O-]
Carboxylic Acid $-COOH$ C(=O)O
Ketone (Carbonyl) $-CO-$ C=O
Amino $-NH_{2}$ N
Cyano $-CN$ C#N
Methyl $CH_{3}-$ C
Isobutyryl $-C(CH_{3})_{2}O$ C(C)=O
Oxime $-C=NOH$ C(=N)O
Methoxy $OCH_{3}$ OC
Amide $-CONH_{2}$ C(N)=O
Ethyl $CH_{2}CH_{3}$ CC
Propyl $CH_{2}CH_{2}CH_{3}$ CCC
Vinyl $CH_{2}=CH-$ C=C
Ethynyl $C\equiv CH$ C#C